### v2

@@ 3.3 An inverse-probability weighted estimator
In this work, we consider IPW-based estimators of population WQTEs (IPW-WQTEs). Specifically, supposing first that the propensity score and double sampling probabilities are known, we consider for any $\tau\in(0,1)$ the estimator defined as approximately solving the following estimating equation:
@@
$E\left(\left\{R+\frac{S(1-R)}{\eta(Z,X)}\right\}W^{g}(Z,X)(1,Z)\left[\tau-\mathbbm{1}\{Y<\beta^{g*}_{0}(\tau)+\beta^{g*}(\tau)Z\}\right]\right)=0$ (6)
at the true values $\{\beta^{g*}_{0}(\tau),\beta^{g*}(\tau)\}$, which motivates our approach. Finally, denote the solution to (5) as $\{\widehat{\beta}_{0}^{g}(\tau),\widehat{\beta}^{g}(\tau)\}$.
In practice, as discussed in Section  3.2, the true propensity score and, possibly, the double-sampling probability, will need to be estimated. Following Assumption 5, suppose the parameters indexing the models $\eta(Z,X;\kappa)$ and $e(X;\gamma)$ are estimated in-sample (i.e., using the data we have), obtained as the solutions to estimating equations based on estimating functions $\phi_{\kappa}$ and $\phi_{\gamma}$, respectively. Plugging in the estimated propensity scores and double-sampling probabilities into expression (5) yields the proposed estimator as approximately solving an equation:
$\mathbbm{P}_{n}\left(\left\{R+\frac{S(1-R)}{\eta(Z,X;\hat{\kappa})}\right\}{W^{g}}(Z,X;\hat{\gamma})(1,Z)\left[\tau-\mathbbm{1}\{Y<\beta^{g}_{0}(\tau)+\beta^{g}(\tau)Z\}\right]\right)=0.$ (7)
Equivalently, letting $\theta=\{\kappa,\gamma,\beta^{g}_{0}(\tau),\beta^{g}(\tau)\}$ and $\Phi_{\theta}=(\phi_{\kappa},\phi_{\gamma},\phi_{\theta})$ with
$\phi_{\theta}=\left\{R+\frac{S(1-R)}{\eta(Z,X;\kappa)}\right\}W^{g}(Z,X;\gamma)(1,Z)\left[\tau-\mathbbm{1}\{Y<\beta^{g}_{0}(\tau)+\beta^{g}(\tau)Z\}\right],$ (8)
@@ 3.4 Asymptotic theory for fixed quantiles
Theorem 1.
Under Assumptions 15, for any fixed $\tau\in(0,1)$,
$n^{1/2}\{\widehat{\beta}^{g}(\tau)-\beta^{g*}(\tau)\}=e_{2}^{\top}\Sigma^{-1}(\tau)\mathbbm{G}_{n}(\psi_{\theta^{*}})+o_{p}(1),$
where $e_{2}=(0,1)^{\top}$,
@@
$\psi_{\theta}=V_{\kappa,3}V^{-1}_{\kappa,1}\phi_{\kappa}+V_{\gamma,3}V^{-1}_{\gamma,2}\phi_{\gamma}-\phi_{\theta},$
$V_{\kappa,1}$ and $V_{\gamma,2}$ respectively are the derivative of $E(\phi_{\kappa})$ and $E(\phi_{\gamma})$, and $V_{\kappa,3}$ and $V_{\gamma,3}$ sequentially are the derivatives of $E(\phi_{\theta})$ at $\kappa$ and $\gamma$, and $\mathbbm{G}_{n}f=\sqrt{n}(\mathbbm{P}_{n}-P)f$.
Note that the vector $e_{2}$ is used to derive the linear representation for the proposed IPW-WQTE estimator $\widehat{\beta}^{g}(\tau)$, which is one of the two coefficients in equation (7). Assumption 4 ensures that the matrix $\Sigma(\tau)$ is invertible. Briefly, the proof of this theorem first involves showing that the class $\{\Phi_{\theta}:||\theta-\theta^{*}||<\delta\}$ is $P$-Donsker for some $\delta>0$ with finite envelope function. Theorem 19.26 of van der Vaart (2000) can then be invoked directly. The full proof of the above theorem is provided in Section S.2. of the Supplementary Materials. Crucially, Theorem 1 establishes that the proposed IPW-WQTE estimator, $\widehat{\beta}^{g}(\tau)$, can be written in a form of linear representation as a normalized sum of i.i.d. random variables, with a negligible remainder, so that, in turn, the central limit theorem can be applied:
Corollary 1.
Suppose that the assumptions in Theorem 1 hold. For any fixed $\tau\in(0,1)$, $n^{1/2}\{\widehat{\beta}^{g}(\tau)-\beta^{g*}(\tau)\}$ converges in distribution to $N(0,e_{2}^{\top}\Sigma^{-1}cov(\psi_{\theta^{*}})\Sigma^{-1}e_{2})$ as $n\rightarrow\infty$.
@@ 3.5 Asymptotic theory for the entire quantile process
For settings where interest lies in estimating causal WQTEs across a range of quantiles, we extend the asymptotic theory regarding $\widehat{\beta}^{g}(\tau)$ presented above to the entire process $\{\widehat{\beta}^{g}(\tau):\tau\in\mathcal{T}\}$ with $\mathcal{T}$ some compact subset of $(0,1)$. For an arbitrary set $\mathcal{T}$, let $\ell^{\infty}(\mathcal{T})$ denote the space of all bounded functions from set $\mathcal{T}$ to $\mathbbm{R}$ equipped with the uniform norm $\|f\|_{\mathcal{T}}=\sup_{\tau\in\mathcal{T}}|f(\tau)|$. Specifically, the following Theorem establishes key asymptotic properties of the process $M_{n}=\left\{n^{1/2}\{\widehat{\beta}^{g}(\tau)-\beta^{g*}(\tau)\}:\tau\in\mathcal{T}\right\}$:
Theorem 2.
Under the conditions of Theorem 1 hold, for any $\tau_{1},\tau_{2}\in\mathcal{T}$, the sequence of maps $M_{n}$ convergences weakly to a measurable tight process $M_{\infty}$ in $\ell^{\infty}(\mathcal{T})$ as $n\rightarrow\infty$, where $M_{\infty}$ is a centered Gaussian process with covariance
$cov\{M_{\infty}(\tau_{1})M_{\infty}(\tau_{2})\}=e^{\top}_{2}\Sigma^{-1}(\tau_{1})\Omega(\tau_{1},\tau_{2})\Sigma^{-1}(\tau_{2})e_{2}$
@@
See the Supplementary Materials for a proof.
@@ 3.6 Variance estimation via the bootstrap
$\widehat{\beta}^{g^{*}}(\tau)\in\mathrm{argmin}_{\beta^{g}_{0}(\tau),\beta^{g}(\tau)}\mathbbm{P}_{n}$ [$\Bigg{[}\left\{R+\frac{S(1-R)}{\eta(Z,X;\hat{\kappa})}\right\}{W^{g}}(Z,X;\hat{\gamma})\rho_{\tau}\{Y-\beta^{g}_{0}(\tau)-\beta^{g}(\tau)Z\}$
@@
where $\rho_{\tau}(u)=\tau\{\tau-\mathbbm{1}(u\leq\tau)\}$ is the check function, the parameters $\kappa,\gamma$ are estimated from the original data, and
$\mathbbm{U}^{*}(\beta^{g}_{0}(\tau),\beta(\tau);\tau)=\ \frac{1}{\sqrt{n}}\sum_{i=1}^{n}\left\{R_{i}+\frac{S_{i}(1-R_{i})}{\eta(Z_{i},X_{i};\hat{\kappa})}\right\}{W^{g}}(Z_{i},X_{i};\hat{\gamma}){\{\beta^{g}_{0}(\tau)+\beta^{g}(\tau)Z_{i}\}}\{\tau-\mathbbm{1}(U_{i}\leq\tau)\},$
where $\{U_{i};i=1,\ldots,n\}$ is a sequence of i.i.d. $U(0,1)$ random variables. The following theorem justifies the validity of this gradient bootstrap estimator of $\beta^{g}(\tau)$ and thereafter its use to construct data-driven uniform confidence bands for $\beta^{g}(\tau),\tau\in\mathcal{T}$.
Theorem 3.
Suppose that the conditions in Theorem 1 hold. Then the sequence of maps $\hat{M}_{n}=\left[n^{1/2}\{\widehat{\beta}^{g}(\tau)-\widehat{\beta}^{g*}(\tau)\},\tau\in\mathcal{T}\right]$ convergences weakly to a measurable tight process $M_{\infty}$ in $\ell^{\infty}(\mathcal{T})$ defined in Theorem 2.
The proof of Theorem 3 can be found in the Supplementary Materials. As mentioned above, Theorem 3 allows us to compute (asymptotically) uniformly valid confidence bands for the process $\beta^{g}(\tau),\tau\in\mathcal{T}$ by simulating the process $\mathbbm{U}^{*}(\beta^{g}_{0}(\tau),\beta(\tau);\tau)$. The algorithm is provided in the Supplementary Materials.
In practice, if researchers do not have a predefined quantile level of interest and aim to estimate WQTEs across multiple quantile levels, Theorems 2 and 3 ensure asymptotically valid inference across an arbitrary compact set in $(0,1)$. Importantly, the uniform confidence bands are useful for making global inference across a compact set $\mathcal{T}$ rather than just a specific point; any null hypothesis that lies outside of the uniform confidence bands, even at a single quantile, can be rejected at a pre-specified significant level. For example, if the band fails to contain 0 at even one quantile in the set $\mathcal{T}$ of interest, we can reject the null hypothesis that there is no causal effect across $\mathcal{T}$.
@@ 4.1 Data generation
For each simulation scenario we generated 10,000 datasets, each with sample size $n=10,000$. For each observation within each dataset, we generated two covariates: $X_{1}\sim U(0,1)$ and $X_{2}\sim U(0,2)$. Letting $X\equiv(X_{1},X_{2})$, a binary ‘treatment’ variable was generated as $Z\mid X\sim\mbox{Bernoulli}\{e(X)\}$, with $e(X)=\{1+\exp(-0.5+0.5X_{1}+0.5X_{2})\}^{-1}$. Finally, the outcome, $Y$, was generated according to the following linear regression model: $Y=1+Z+X_{1}+X_{2}+(1+\rho Z)\varepsilon$, where $\varepsilon$ is the error term and $\rho$ is a constant controlling the heteroscedasticity of QTEs across quantiles $\tau\in(0,1)$. For the error terms, we considered a Pareto distribution with location parameter 1 and shape parameter 5. With this choice, the induced marginal distribution of outcome $Y$ is highly right skewed, as shown in Figure S.1 in the Supplementary Material. For $\rho$, we considered two scenarios: (i) $\rho=0$, so that the QTE is homogeneous in across quantile levels; and, (ii) $\rho=1.5$, so that the QTEs are heterogeneous across quantile levels. The true QTE is 1 for the homogeneous setting, and the true QTEs vary from 2.67 to 3.10 for the heterogeneous setting; see Figure S.2 in the Supplementary Material.
@@
Finally, we considered double-sampling of those with $R=0$ within strata defined jointly by $(Z,X_{1},X_{2})$, with $X_{1}$ and $X_{2}$ dichotomized using cut-off points of 0.5 and 1, respectively. We pre-defined the stratum-specific sample size varying from 0 and 240 in each of the eight strata. The specific sample size for each stratum is listed in Section S4 of the Supplementary Materials. To ensure the non-informative double-sampling assumption, a stratum-specific sample was randomly ‘selected’ in each stratum. This procedure ‘selected’ a total of double-sampled 800 individuals with initially missing outcomes. Overall these choices resulted in double-sampling of approximately 22% of individuals with original missing data.
@@ 4.2 Analysis
We consider estimation of the population QTE, $\beta^{g}(\tau)$ with $g(X)=1$, with values of $\tau$ ranging from 0.1 to 0.9 with a grid spacing of 0.1. Specifically, we considered four estimators: (I) an estimator based on the full data, with $e(X)$ estimated via maximum likelihood; (II) an estimator based on the initial incomplete data alone (i.e., only including those with complete outcome data at the outset and not including the double-sampled data), with $e(X)$ estimated via maximum likelihood; (III) an estimator based on the initial incomplete data coupled with the double-sampled data, with $e(X)$ set at their true values and $\eta(Z,X)$ estimated via maximum likelihood; and, (IV) an estimator based on the initial incomplete data coupled with the double-sampled data, with both $e(X)$ and $\eta(Z,X)$ estimated via maximum likelihood. Note, throughout, estimates of $e(X)$ were based on a correctly-specific logistic regression. Furthermore, since we performed a stratified sampling with fixed sample sizes, and the eight potential strata had random sample sizes (due to the re-generation of the covariate data), there is no single set of ‘true’ values for the double-sampling probabilities. We thus estimated $\eta(Z,X)$ via a saturated logistic regression on the eight strata.
Asymptotic variance calculations for proposed estimators (III) and (IV) were computed assuming the true underlying distribution family. Specifically, for estimator (III), the function $\psi_{\theta}$ from Theorem 1 simplifies to $V_{\kappa,3}V^{-1}_{\kappa,1}\phi_{\kappa}-\phi_{\theta}$. The asymptotic variance was then calculated with the true $e(x)$ and estimated parameters $\{\widehat{\kappa},\widehat{\beta}_{0}^{g}(\tau),\widehat{\beta}^{g}(\tau)\}$ plugged into $e_{2}^{\top}\Sigma^{-1}cov(\psi_{\theta})\Sigma^{-1}e_{2}$. For estimator (IV), we first estimated parameters $\widehat{\theta}=\{\widehat{\kappa},\widehat{\gamma},\widehat{\beta}_{0}^{g}(\tau),\widehat{\beta}^{g}(\tau)\}$, then plugged them into the corresponding asymptotic variance expression of (IV), as presented in Theorem 1 and Corollary 1. Additionally, pointwise bootstrap variance estimates were computed for estimators (III) and (IV), for each $\tau$ considered. Finally, we constructed 95% confidence bands using the gradient bootstrap method described in Section 3.5. To assess coverage for the latter, we determined whether a given confidence band contained the true QTE value for each quantile level in the sequence considered. If the band contained all of the true values, we consider it successful coverage for the given realization; otherwise, it is considered a failure. The empirical coverage probability is the proportion of realizations with successful coverage of the true parameters.

### v3

@@ 3.3 An inverse-probability weighted estimator
In this work, we consider IPW-based estimators of population WQTEs (IPW-WQTEs). Specifically, supposing first that the propensity score and double sampling probabilities are known, we consider for any $\tau\in(0,1)$ the estimator defined as approximately solving:
@@
$E\left(\left\{R+\frac{S(1-R)}{\eta(Z,X)}\right\}W^{g}(Z,X)\begin{pmatrix}1\\[-5.0pt] Z\end{pmatrix}\left[\mathbbm{1}\{Y<\beta^{g*}_{0}(\tau)+\beta^{g*}(\tau)Z\}\right]-\tau\right)=0$ (5)
at the true values $\{\beta^{g*}_{0}(\tau),\beta^{g*}(\tau)\}$, which motivates our approach.
In practice, as discussed in Section 3.2, the true propensity score and, possibly, the double-sampling probability, will need to be estimated. Following Assumption 3.2, suppose the parameters indexing the models $\eta(Z,X;\kappa)$ and $e(X;\gamma)$ are estimated in-sample (i.e., using the data we have), obtained as the solutions to estimating equations based on estimating functions $\phi_{\kappa}$ and $\phi_{\gamma}$, respectively. For example, $\eta(Z,X;\kappa)$ and $e(X;\gamma)$ can be estimated using logistic regressions, with $\phi_{\kappa}$ and $\phi_{\gamma}$ representing the corresponding score functions. An example of such a score function is $\phi_{\gamma}=\sum_{i=1}^{n}[y_{i}-\{1+\exp(-\boldsymbol{x}_{i}^{\top}\gamma)\}^{-1}]\boldsymbol{x}_{i}^{\top}$. Subsequently, $\widehat{\kappa}$ and $\widehat{\gamma}$ denote the maximum likelihood estimators for these parameters.
Plugging in the estimated propensity scores and double-sampling probabilities into expression (4) yields the proposed estimator as approximately solving an equation:
$\mathbbm{P}_{n}\left(\left\{R+\frac{S(1-R)}{\eta(Z,X;\widehat{\kappa})}\right\}{W^{g}}(Z,X;\widehat{\gamma})\begin{pmatrix}1\\[-5.0pt] Z\end{pmatrix}\left[\mathbbm{1}\{Y<\beta^{g}_{0}(\tau)+\beta^{g}(\tau)Z\}-\tau\right]\right)=0.$ (6)
Equivalently, letting $\theta=\{\kappa,\gamma,\beta^{g}_{0}(\tau),\beta^{g}(\tau)\}$ and $\Phi_{\theta}=(\phi_{\kappa},\phi_{\gamma},\phi_{\beta})$ with
$\phi_{\beta}=\left\{R+\frac{S(1-R)}{\eta(Z,X;\kappa)}\right\}W^{g}(Z,X;\gamma)\begin{pmatrix}1\\[-5.0pt] Z\end{pmatrix}\left[\mathbbm{1}\{Y<\beta^{g}_{0}(\tau)+\beta^{g}(\tau)Z\}-\tau\right],$ (7)
@@ 3.4 Asymptotic theory for fixed quantiles
Theorem 3.1
[…]
$n^{1/2}\{\widehat{\beta}^{g}(\tau)-\beta^{g*}(\tau)\}=\boldsymbol{e}_{2}^{\top}\Sigma^{-1}(\tau)\mathbbm{G}_{n}(\psi_{\theta^{*}})+o_{p}(1),$
where $\boldsymbol{e}_{2}=(0,1)^{\top}$,
@@
$\psi_{\theta}=V_{\kappa,3}V^{-1}_{\kappa,1}\phi_{\kappa}+V_{\gamma,3}V^{-1}_{\gamma,2}\phi_{\gamma}-\phi_{\beta},$
$V_{\kappa,1}$ and $V_{\gamma,2}$ respectively are the derivative of $E(\phi_{\kappa})$ and $E(\phi_{\gamma})$, and $V_{\kappa,3}$ and $V_{\gamma,3}$ sequentially are the derivatives of $E(\phi_{\beta})$ at $\kappa$ and $\gamma$, and $\mathbbm{G}_{n}f=\sqrt{n}(\mathbbm{P}_{n}-P)f$.
While complete details are given in Section S.2. of the Supplementary Materials, briefly, the proof of Theorem 3.1, which includes proof of consistency of $\widehat{\beta}^{g}(\tau)$, first involves showing that the class $\{\Phi_{\theta}:||\theta-\theta^{*}||<\delta\}$ is $P$-Donsker for some $\delta>0$ with finite envelope function. Theorem 19.26 of van der Vaart (2000) can then be invoked directly. Note that the vector $\boldsymbol{e}_{2}$ is used to derive the linear representation for the proposed IPW-WQTE estimator $\widehat{\beta}^{g}(\tau)$, which is one of the two coefficients in equation (6). Crucially, Theorem 3.1 establishes that the proposed IPW-WQTE estimator, $\widehat{\beta}^{g}(\tau)$, can be written in a form of linear representation as a normalized sum of i.i.d. random variables, with a negligible remainder, so that, in turn, the central limit theorem can be applied. Remarkably, the variance arising from estimating the nuisance parameters $\kappa$, $\gamma$ and $\beta_{0}^{g}(\tau)$ is encompassed withing $\psi_{\theta}$.
Corollary 3.2
Assume the assumptions in Theorem 3.1 hold. For any fixed $\tau\in(0,1)$, $n^{1/2}\{\widehat{\beta}^{g}(\tau)-\beta^{g*}(\tau)\}$ converges in distribution to $N(0,\boldsymbol{e}_{2}^{\top}\Sigma^{-1}cov(\psi_{\theta^{*}})\Sigma^{-1}\boldsymbol{e}_{2})$ as $n\rightarrow\infty$.
@@ 3.5 Asymptotic theory for the entire quantile process
For settings where interest lies in estimating causal WQTEs across a range of quantiles, we extend the asymptotic theory regarding $\widehat{\beta}^{g}(\tau)$ presented above to the entire process $\{\widehat{\beta}^{g}(\tau):\tau\in\mathcal{T}\}$ with $\mathcal{T}$ some compact subset of $(0,1)$. For an arbitrary set $\mathcal{T}$, let $\ell^{\infty}(\mathcal{T})$ denote the space of all bounded functions from set $\mathcal{T}$ to $\mathbbm{R}$ equipped with the uniform norm $\|f\|_{\mathcal{T}}=\sup_{\tau\in\mathcal{T}}|f(\tau)|$. Specifically, the following Theorem establishes key asymptotic properties of the process $M_{n}=\left[n^{1/2}\{\widehat{\beta}^{g}(\tau)-\beta^{g*}(\tau)\}:\tau\in\mathcal{T}\right]$:
Theorem 3.3
[…]
$cov\{M_{\infty}(\tau_{1})M_{\infty}(\tau_{2})\}=\boldsymbol{e}^{\top}_{2}\Sigma^{-1}(\tau_{1})\Omega(\tau_{1},\tau_{2})\Sigma^{-1}(\tau_{2})\boldsymbol{e}_{2}$
@@
The proof of Theorem 3.3 can be found in Section S.2 of the Supplementary Materials.
@@ 3.6 Variance estimation via the bootstrap
$\widehat{\beta}^{g^{*}}(\tau)\in\mathrm{argmin}_{\beta^{g}_{0}(\tau),\beta^{g}(\tau)}\mathbbm{P}_{n}$ [$\Bigg{[}\left\{R_{i}+\frac{S_{i}(1-R_{i})}{\eta(Z_{i},X_{i};\widetilde{\kappa})}\right\}{W^{g}}(Z_{i},X_{i};\widetilde{\gamma})\rho_{\tau}\{Y_{i}-\beta^{g}_{0}(\tau)-\beta^{g}(\tau)Z_{i}\}$
@@
where $\rho_{\tau}(u)=\tau\{\tau-\mathbbm{1}(u\leq\tau)\}$ is the check function, $D=\{(R_{i},S_{i},(R_{i}+S_{i})Y_{i},Z_{i},X_{i}):1\leq i\leq n\}$ is the original data, $\widetilde{\kappa}$ and $\widetilde{\gamma}$ are the estimated nuisance parameters using the bootstrap sample $D^{b}=\{(R^{b}_{i},S^{b}_{i},(R^{b}_{i}+S^{b}_{i})Y^{b}_{i},Z^{b}_{i},X^{b}_{i}):1\leq i\leq n\}$, and
$\mathbbm{U}^{*}(\beta^{g}_{0}(\tau),\beta(\tau);\tau)=\ \frac{1}{\sqrt{n}}\sum_{i=1}^{n}\left\{R_{i}+\frac{S_{i}(1-R_{i})}{\eta(Z_{i},X_{i};\widetilde{\kappa}})\right\}{W^{g}}(Z_{i},X_{i};\widetilde{\gamma}){\{\beta^{g}_{0}(\tau)+\beta^{g}(\tau)Z_{i}\}}\{\tau-\mathbbm{1}(U^{b}_{i}\leq\tau)\},$
where for each $D^{b}$, $\{U^{b}_{i}:1\leq i\leq n\}$ is a sequence of i.i.d. $U(0,1)$ random variables. The following theorem justifies the validity of this gradient bootstrap estimator of $\beta^{g}(\tau)$ and thereafter its use to construct data-driven uniform confidence bands for $\beta^{g}(\tau),\tau\in\mathcal{T}$.
Theorem 3.4
Suppose that the conditions in Theorem 3.3 hold. Then the sequence of maps $\widehat{M}_{n}=\left[n^{1/2}\{\widehat{\beta}^{g}(\tau)-\widehat{\beta}^{g*}(\tau)\}:\tau\in\mathcal{T}\right]$ convergences weakly to a measurable tight process $M_{\infty}$ in $\ell^{\infty}(\mathcal{T})$ defined in Theorem 3.3.
The proof of Theorem 3.4 can be found in Section S.2 of the Supplementary Materials. As mentioned above, Theorem 3.4 allows us to compute (asymptotically) uniformly valid confidence bands for the process $\beta^{g}(\tau),\tau\in\mathcal{T}$ by simulating the process $\mathbbm{U}^{*}(\beta^{g}_{0}(\tau),\beta(\tau);\tau)$. The algorithm is provided in Section S.3 of the Supplementary Materials.
In practice, if researchers do not have a predefined quantile level of interest and aim to estimate WQTEs across multiple quantiles, Theorems 3.3 and 3.4 ensure asymptotically valid inference over any compact set in $(0,1)$. Importantly, the uniform confidence bands facilitate global inference across a compact set $\mathcal{T}$ rather than just a specific point; any null hypothesis that lies outside of the uniform confidence bands, even at a single quantile, can be rejected at a pre-specified significant level. For example, if the band fails to contain 0 at any quantile in $\mathcal{T}$, we can reject the null hypothesis that there is no causal effect across $\mathcal{T}$.
@@ 4.1 Data generation
For each simulation scenario we generated 10,000 datasets, each with sample size $n=10,000$. For each observation within each dataset, we generated two covariates: $X_{1}\sim U(0,1)$ and $X_{2}\sim U(0,2)$. Letting $X\equiv(X_{1},X_{2})$, a binary ‘treatment’ variable was generated as $Z\mid X\sim\mbox{Bernoulli}\{e(X)\}$, with $e(X)=\{1+\exp(-0.5+0.5X_{1}+0.5X_{2})\}^{-1}$. Finally, the outcome $Y$ was generated using the linear model $Y=1+Z+X_{1}+X_{2}+(1+\rho Z)\varepsilon$, where $\varepsilon$ is the error term and $\rho$ is a constant controlling the heteroscedasticity of QTEs across quantiles $\tau\in(0,1)$. We used a Pareto distribution for $\varepsilon$, inducing a right-skewed marginal distribution of $Y$ (Figure S.1). We considered two cases for $\rho$: (i) $\rho=0$, resulting in homogeneous QTEs of 1; and (ii) $\rho=1.5$, resulting in heterogeneous QTEs from 2.67 to 3.10 (Figure S.2).
@@
Finally, we considered double-sampling those with $R=0$ within strata defined by $(Z,X_{1},X_{2})$, where $X_{1}$ and $X_{2}$ were dichotomized at 0.5 and 1, respectively. We pre-defined the stratum-specific sample size varying from 0 and 240 in each of the eight strata (see Table S.1 and S.2 in Section S.4.1 of the Supplementary Materials). To ensure the non-informative double-sampling assumption, a stratum-specific sample was randomly ‘selected’ in each stratum. This procedure ‘selected’ a total of double-sampled 800 individuals with initially missing outcomes. Overall these choices resulted in double-sampling of approximately 22% of individuals with original missing data.
@@ 4.2 Analysis
We estimate the population QTE, $\beta^{g}(\tau)$ with $g(X)=1$, for $\tau$ from 0.1 to 0.9 with increments of 0.1. Specifically, we considered five estimators: (I) an estimator based on the full data, with $e(X)$ estimated via maximum likelihood; (II) an estimator based on the initial incomplete data alone (excluding double-sampled data), with $e(X)$ estimated via maximum likelihood; (III) an estimator based on the initial incomplete data coupled with the double-sampled data, with $e(X)$ set at their true values and $\eta(Z,X)$ estimated via maximum likelihood; (IV) an estimator based on the initial incomplete data coupled with the double-sampled data, with both $e(X)$ and $\eta(Z,X)$ estimated via maximum likelihood; and, (V) an estimator based on the initial incomplete data assuming MAR, with both $e(X)$ and $\pi(Z,X):=pr(R=1\mid Z,X)$ estimated via maximum likelihood. Specifically, estimator (V) approximately solves an equation $\mathbbm{P}_{n}\left(\frac{R}{\pi(Z,X;\widehat{\alpha})}{W^{g}}(Z,X;\widehat{\gamma})(1,Z)^{\top}\left[\mathbbm{1}\{Y<\beta^{g}_{0}(\tau)+\beta^{g}(\tau)Z\}-\tau\right]\right)=0$. Note, throughout, estimates of $e(X)$ and $\eta(Z,X)$ were based on a correctly-specific logistic regression. Furthermore, since we performed a stratified sampling with fixed sample sizes, and the eight potential strata had random sample sizes (due to the re-generation of the covariate data), there is no single set of ‘true’ values for the double-sampling probabilities. We thus estimated $\eta(Z,X)$ via a saturated logistic regression on the eight strata.
Asymptotic variance calculations for proposed estimators (III) and (IV) were computed assuming the true underlying distribution family. Specifically, for estimator (III), the function $\psi_{\theta}$ from Theorem 3.1 simplifies to $V_{\kappa,3}V^{-1}_{\kappa,1}\phi_{\kappa}-\phi_{\beta}$. The asymptotic variance was then calculated with the true $e(x)$ and estimated parameters $\{\widehat{\kappa},\widehat{\beta}_{0}^{g}(\tau),\widehat{\beta}^{g}(\tau)\}$ plugged into $\boldsymbol{e}_{2}^{\top}\Sigma^{-1}cov(\psi_{\theta})\Sigma^{-1}\boldsymbol{e}_{2}$. For estimator (IV), we first estimated parameters $\widehat{\theta}=\{\widehat{\kappa},\widehat{\gamma},\widehat{\beta}_{0}^{g}(\tau),\widehat{\beta}^{g}(\tau)\}$, then plugged them into the corresponding asymptotic variance expression of (IV), as presented in Theorem 3.1 and Corollary 3.2. Pointwise bootstrap variance estimates were computed for estimators (III) and (IV) at each $\tau$ considered, and the 95% confidence bands were constructed using the gradient bootstrap method (Section 3.5). Coverage was assessed by determining if the band included the true QTE for each quantile; if all values were included, coverage was successful for the given realization; otherwise, it is considered a failure. The empirical coverage probability is the proportion of realizations with successful coverage of the true parameters.